. Total revenue from the sale of X is given by the equation R=60Q-Q2. Calculate the value of marginal revenue when the point price elasticity of demand when marginal revenue is -2.



1
Expert's answer
2021-04-11T20:02:40-0400

If total revenue "R = 60Q - Q^2," then marginal revenue function is "MR = R'(Q) = 60 - 2Q."

The demand function is:

P = R/Q = 60 - Q.

Ed = -1×P/Q = -2,

P/Q = 2,

P = 2Q,

60 - Q = 2Q,

Q = 20 units,

MR = 60 - 2×20 = 20.